### v48

# 0.5 eV QCD Axion Cosmology

Abstract
A simple yet compelling physical picture is proposed for the nature of dark matter, dark energy, and the Big Bang. The proposal leads to a prediction with 0.53 parts per million precision for the 0.5 eV rest-mass energy of the quantum chromodynamic (QCD) axions that form the dark matter within this picture. A straightforward resolution of the Hubble tension follows within this picture from the agreement between the prediction for the present expansion rate of the universe and the best present determination from observations of the late universe.
[…]
show]nbrayali@gmail.com
1 Introduction
In 2022 the Hubble tension between late-universe observations of the present expansion rate (A. G. Riess et al., 2022) and predictions within the standard cold dark matter cosmology with cosmological constant ($\rm\Lambda CDM$}) based on early-universe observations (Planck Collaboration et al., 2020; S. Aiola et al., 2020) reached the conventional $5.0\sigma$ threshold for the discovery of a new phenomenon in astrophysics. Several years before this conventional threshold was reached, astrophysicists were already stimulated by the accumulating evidence for the existence of a significant tension, and they began suggesting innovative alternative cosmologies that might reconcile early- and late-universe observations (V. Poulin et al., 2019). In 2024 the tension was confirmed on the late-universe side with James Webb Space Telescope (JWST) observations that rejected unrecognized crowding of Cepheid variable star photometry as an explanation of the Hubble tension at $8\sigma$ confidence (A. G. Riess et al., 2024a), and that validated the Hubble Space Telescope (HST) distance measurements on which the tension rests (A. G. Riess et al., 2024b).
On the early-universe side, the Dark Energy Spectroscopic Instrument (DESI) first-year observations of the peak in the galaxy-galaxy two-point correlation function confirmed the Hubble tension in 2025, provided the results are interpreted within the standard $\Lambda{\rm CDM}$ cosmology (A. G. Adame et al., 2025). Similarly, in 2023 an independent combination of ground-based and space-based observations of the cosmic microwave background (CMB) power spectrum of angular fluctuations confirmed the tension, again within the framework of $\Lambda{\rm CDM}$ (L. Balkenhol et al., 2023). In 2024 the first comprehensive book-length review concluded that the Hubble tension remains an unsolved mystery (A. G. Riess, 2024).
The purpose of this Letter is to propose a simple yet compelling physical picture for the nature of dark matter, dark energy, and the Big Bang. However, the picture offers a straightforward resolution of the Hubble tension, and this resolution is, therefore, the primary astrophysical motivation for considering the picture, despite the fact that, historically, the picture emerged from quite different motivations and considerations (See Section 2). In brief, the picture leads to the following high-precision prediction from first-principles for the value of the present expansion rate of universe (See Section 3 for details):
$H_{0}=71.939~5~(76)~{\rm km\cdot sec^{-1}\cdot Mpc^{-1}}~[{\rm 0.11~per~mille}].$ (1)
The predicted expansion rate in Eq. (1) roughly agrees, within the percent-level uncertainty of the comparison, with the best present late-universe determination using the relation between redshift and distance for Type Ia supernovae by the SH0ES collaboration (A. G. Riess et al., 2025)
$H_{0}({\rm SH0ES})=73.18~(88)~{\rm km\cdot sec^{-1}\cdot Mpc^{-1}}~[{\rm 12~per~mille}].$ (2)
Correspondingly, the predicted expansion rate in Eq. (1) differs sharply, by $10\sigma$, from the best present $\Lambda{\rm CDM}$ prediction using early-universe observations (Planck Collaboration et al., 2019, 2020)
$H_{0}(\Lambda{\rm CDM})=67.9~(4)~{\rm km\cdot sec^{-1}\cdot Mpc^{-1}}~[{\rm 5.9~per~mille}].$ (3)
Within the physical picture presented in this Letter, the straightforward resolution of the Hubble tension is simply that the late-universe observations are basically correct (on this point, see also D. Scolnic et al. (2025)), as are the $\Lambda{\rm CDM}$ predictions for the value of the cosmological constant and for the value of the present energy density of dark matter in the universe, but that the value of the present energy density of baryons is more than double the value predicted by $\Lambda{\rm CDM}$ (See Section 3 for the actual values of the corresponding cosmological parameters and for the comparison with $\Lambda{\rm CDM}$).
In a spatially flat universe, the value of the present expansion rate, $H_{0}=100h~{\rm km~sec^{-1}~Mpc^{-1}}$, is linked to the values of the cosmological parameters of the cosmological constant, $\Lambda$, the dark matter energy density, $u_{A}$, the baryon energy density, $u_{B}$, and the photon energy density, $u_{\gamma}$, by the simple relation (P. J. E. Peebles, 1993, pg. 100)
$h^{2}=\Omega_{A}h^{2}+\Omega_{B}h^{2}+\Omega_{\gamma}h^{2}+\Omega_{\Lambda}h^{2},$ (4)
where the cosmological parameter is $\Omega_{X}h^{2}=(8\pi/3)(Gu_{X}/c^{2})(h/H_{0})^{2}$ for energy density $u_{X}$, with $X=A,B,\gamma$, and it is $\Omega_{\Lambda}h^{2}=(c^{2}\Lambda/3)(h/H_{0})^{2}$ for the cosmological constant $\Lambda$. To resolve the Hubble tension, we propose a new cosmology, based on dark matter in the form of fundamental particles known as axions (S. Weinberg, 1978; F. Wilczek, 1978), that leads to predictions for the values of the cosmological parameters, $\Omega_{A}h^{2}$ and $\Omega_{B}h^{2}$, which have the same precision as the cosmological parameter for photons, $\Omega_{\gamma}h^{2}=2.473~(2)\times 10^{-5}~[{\rm 0.88~per~mille}]$, set by the precision of the best present estimate of the CMB temperature $T_{\gamma}=2.725~5~(6)~{\rm K}~[{\rm 0.22~per~mille}]$ (D. J. Fixsen, 2009). Together with the first-principles prediction, within this axion cosmology, for the value of $\Omega_{\Lambda}h^{2}$, to an astounding precision of 22 parts per million (ppm), the prediction for $H_{0}$ given in Eq. (1) is obtained using Eq. (4).
The linchpin of the analysis leading to this straightforward resolution of the Hubble tension is the following remarkable, numerically exact relation for the ratio of the energy density of baryons to the energy density of axions in the present universe (See Section 4 for the derivation)
$\frac{\Omega_{B}h^{2}}{\Omega_{A}h^{2}}=\frac{2}{\pi^{2}}\left(x_{k}y_{T}\right)^{2}\left(\frac{1+\delta}{1+\epsilon}\right)=0.393~565~531~\left(\frac{1+\delta}{1+\epsilon}\right),$ (5)
where the small dimensionless quantity, $\epsilon=(4m_{p}/m_{\alpha}-1)Y_{P}\approx 0.002$, corrects for Big Bang nucleosynthesis of primordial helium with nucleus mass, $m_{\alpha}$, and primordial abundance by mass, $Y_{P}$, and the relativistic correction is $\delta\approx 0.06$ (See Appendix C), while the numerically exact constant, $x_{k}=2.821~439~372$ (P. J. E. Peebles, 1993, pg. 159), gives the kinetic energy, $\check{E}=x_{k}kT_{\pi}$, for the axions on the peak of the black-body distribution at the temperature of the Big Bang, $T_{\pi}=y_{T}\left(\chi_{\rm QCD}\right)^{1/4}/k$, and the numerically exact constant, $y_{T}=(5/84)^{1/4}=0.493~938~274$, relates this temperature to the topological susceptibility of the quantum chromodynamic (QCD) vacuum, $\chi_{\rm QCD}=(m_{A}c^{2})^{2}f_{A}^{2}$ (S. Weinberg, 1978; F. Wilczek, 1978), given by the axion rest-mass energy, $m_{A}c^{2}$, and the axion decay constant, $f_{A}$. The value of the corresponding ratio within $\Lambda{\rm CDM}$ cosmology is roughly half as large as the value given in Eq. (5) (Planck Collaboration et al., 2019, 2020). This sharp difference resolves the Hubble tension, as shown in Eq. (1), and has implications for a wide range of astrophysical phenomena, including the formation of large-scale structure during the cross-over from radiation-domination to matter-domination (See Section 5), the synthesis of the nuclei of the chemical elements with low atomic mass during the Big Bang (See Section 6), and the spectrum of initial fluctuations in energy density at the Big Bang itself (See Section 7).
A key prediction of the physical picture for dark matter, dark energy, and the Big Bang presented in this Letter is the value of the rest-mass energy of the 0.5 eV QCD axion dark matter to 0.53 parts per million (ppm) precision (See Appendix B for the derivation):
$m_{A}c^{2}=0.504~321~77~(27)~{\rm eV}~[{\rm 0.53~ppm}],$ (6)
This combines with the percent-level precise prediction for the strength of the axion-photon coupling (See Appendix E),
$g_{A\gamma\gamma}=0.68~(2)\times 10^{-10}~{\rm GeV}^{-1}~[{\rm 3~per~cent}].$ (7)
Together these high-precision, first-principles predictions for the key physical parameters of axion dark matter provide, for the first time, the necessary inputs for the analysis of the astrophysical effects of axion dark matter.
Finally, a wide range of astrophysical observations have been used over the past five decades to “constrain” or even “exclude” axion dark matter with the rest-mass energy given in Eq. (6) and the axion-photon coupling given in Eq. (7) (See Section 8). Most recently, in 2025, blank-sky observations with the JWST near-infrared spectrometer NIRSpec and mid-infrared spectrometer MIRI (R. Janish & E. Pinetti, 2025; S. Roy et al., 2025) were used to derive strong limits on the axion-photon coupling constant, $g_{A\gamma\gamma}$, in the mass range 0.1 to 4 eV, that “exclude,” with much greater than 95% confidence, these predicted values (E. Pinetti, 2025). Yet, more than a decade ago, in 2014, observations of horizontal branch stars in globular clusters, combined with stellar evolution that includes the core cooling effects from axion emission due to photon-to-axion conversion in the helium-burning cores of these stars, had already been used to derive a strong “upper bound” on the value of the axion-photon coupling, $g_{A\gamma\gamma}<0.66\times 10^{-10}~{\rm GeV}^{-1}$, independent of the value of $m_{A}$, which “excludes,” with just over 95% confidence, the predicted value in Eq. (7) (A. Ayala et al., 2014).
In light of the apparent success of 0.5 eV QCD axion dark matter in resolving the Hubble tension, these astrophysical “constraints” and “exclusions” raise important questions about the conventional assumptions for a wide range of astrophysical phenomena. From the velocity dispersion of dark matter in the “halo” of the galaxy to the core helium abundance at formation for stars in globular clusters, these astrophysical assumptions underlie the analyses of the effects of axions on observations and simulations. Turning these analyses on their head, we conclude that the physical picture, presented in this Letter, stimulates a new look at these conventional assumptions, which cut across a wide swath of astrophysical phenomena from star formation to supernovae, from galactic structure to galactic chemical evolution, and from the structure to the evolution of pulsars, white dwarfs, red giants, and horizontal branch stars (See Section 9).
2 0.5 eV QCD Axion Physical Picture
The physical picture for dark matter, dark energy, and the Big Bang begins with an observation by Y. B. Zel’Dovich (1967). Stimulated by some of the first quasar absorption line spectroscopy (G. Burbidge, 1967), a number of astrophysicists suggested a value for the cosmological constant $\Lambda\approx+5\times 10^{-56}~{\rm cm}^{-2}$ (N. Kardashev, 1967; J. Shklovsky, 1967; V. Petrosian et al., 1967). In this context, Y. B. Zel’Dovich (1967) observed that the gravitational binding energy, $U_{\Lambda}$, generated by a massless particle and antiparticle, each with the kinetic energy, $E$, separated by the distance, $R=\hbar c/E$,
$U_{\Lambda}$ $=$ $N_{U}G\frac{(E/c^{2})^{2}}{R}$ (8)
$=$ $N_{U}\frac{E^{3}}{E_{\rm Pl}^{2}},$
with $N_{U}$ a dimensionless constant of order $\mathcal{O}(10^{-1})$, and having particle-antiparticle pair number density,
$n_{\Lambda}$ $=$ $N_{n}\frac{1}{R^{3}}$ (9)
$=$ $N_{n}\left(\frac{E}{\hbar c}\right)^{3},$
with the dimensionless constant, $N_{n}$, of order $\mathcal{O}(10^{-1})$ to prevent Bose-Einstein condensation of the particle-antiparticle pairs, gives cosmological “dark energy” with energy density, $u_{\Lambda}=U_{\Lambda}n_{\Lambda}$, of the right order of magnitude to explain the value of the cosmological constant, $\Lambda=8\pi Gu_{\Lambda}/c^{4}$,
$u_{\Lambda}$ $=$ $N_{U}N_{n}\left(\frac{E^{3}}{E_{\rm Pl}^{2}}\right)\left(\frac{E}{\hbar c}\right)^{3}$ (10)
$=$ $1.874~579~(41)\times 10^{-8}~{\rm erg~cm^{-3}}~[{\rm 22~ppm}]$
$\times$ $\left(\frac{N_{U}}{0.1}\right)\left(\frac{N_{n}}{0.1}\right)\left(\frac{E}{105~{\rm MeV}}\right)^{6},$
provided that the kinetic energy, $E$, is of the order of $\mathcal{O}(m_{p}c^{2}\times 10^{-1})$, where $m_{p}c^{2}\approx 940~{\rm MeV}$ is the rest-mass energy of the proton and $E_{\rm Pl}=\sqrt{\hbar c^{5}/G}=1.220~932~(13)\times 10^{22}~{\rm MeV}~[11~{\rm ppm}]$ is the Planck energy scale:
$E$ $=$ $\left(\Omega_{\Lambda}h^{2}\right)^{1/6}\left(\frac{3}{8\pi N_{n}N_{U}}\right)^{1/6}\left(E_{\rm Pl}\right)^{2/3}\left(\hbar\frac{H_{0}}{h}\right)^{1/3}$ (11)
$=$ $86.6~210~(6)~{\rm MeV}~\left(\frac{\Omega_{\Lambda}h^{2}}{0.35}\right)^{1/6}\left(\frac{0.1}{N_{n}}\right)^{1/6}\left(\frac{0.1}{N_{U}}\right)^{1/6}~[7.3~{\rm ppm}].$
In this historical context, the key insight is simply to regard the particle-antiparticle pairs, which are responsible for the dark energy described by the cosmological constant, as forming the dark matter in the universe. The Big Bang, then, is just the moment when massless matter and antimatter bind to form the dark matter, in the form of these particle-antiparticle pairs. As the universe expands, the number density of dark matter, $n_{A}(t)\propto 1/a(t)^{3}$, dilutes with the volume of the expanding universe given by the scale factor of the universe, $a(t)$, at time $t$ after the Bang, and, thus, the kinetic energy of the massless matter and antimatter inside the dark matter, $E(t)\propto a(t)$, must grow to keep constant in time the value of the cosmological constant, $\Lambda\propto n_{A}(t)E(t)^{3}$.
Starting from the value $\check{E}$, given by Eq. (11), this kinetic energy reaches its peak value, $\hat{E}$, at future infinity when the dark matter dissolves and the universe begins contracting. For much of the past four decades, astrophysicists have looked to the Big Bang as a kind of cosmic accelerator for reaching collision energies sufficient to probe phenomena beyond the standard model of particle physics and the standard cold dark matter cosmology (L. Amendola et al., 2018; X. Chen et al., 2017; L. M. Lederman & D. N. Schramm, 1989, pg. 161). Yet, it is a simple historical fact that the last generation of the “classic” particle accelerators, known as synchrocyclotrons and built until the 1950s, already reached beam energies sufficient to make pairs of muons (L. M. Lederman & D. N. Schramm, 1989, pgs. 93–94), and, thus these classic accelerators already achieved, seven decades ago, collision energies greater than those typically experienced in the Big Bang, where the typical kinetic energy of radiation, $\check{E}$, was on the order of, but smaller than, the rest-mass energy of the muon $m_{\mu}c^{2}\approx 106~{\rm MeV}$ (See Eq. (11)).
Instead, the moment in the life-cycle of the universe where the kinetic energy peaks is at future infinity with the value $\hat{E}$. A natural speculation for this maximum particle kinetic energy value, $\hat{E}\sim 10^{18}~{\rm MeV}$, emerges from analysis of the energy dependence of the strengths of the fundamental forces within the standard model of particle physics: It is the energy where these strengths (almost) converge on a “unified” value (L. M. Lederman & D. N. Schramm, 1989, pg. 160). But, this appealing (and enduring) speculation then implies that the values for the kinetic energy of the massless particles and antiparticles inside dark matter at the Big Bang, $\check{E}$, and at future infinity, $\hat{E}$, satisfy a simple and appealing “see-saw” relation, together with the decay constant, $f_{A}=\sqrt{\chi_{\rm QCD}}/(m_{A}c^{2})\sim 10^{10}~{\rm MeV}$, for the 0.5 eV QCD axion with rest-mass energy, $m_{A}c^{2}\sim 0.5~{\rm eV}=5\times 10^{-7}~{\rm MeV}$ (See Eq. (6)):
$\check{E}\hat{E}=f_{A}^{2},$ (12)
where the best present estimate for the topological susceptibility of the QCD vacuum from Chiral Perturbation Theory (CPT) is $\sqrt{\chi_{\rm QCD}}=5~690~(50)~{\rm MeV^{2}}~[{\rm 8.8~per~mille}]$ (M. Gorghetto & G. Villadoro, 2019).
More rigorously, a per mille precise prediction for the value of the kinetic energy at future infinity, $\hat{E}$, emerges from simply identifying the value of the axion decay constant, $f_{A}=\sqrt{\chi_{\rm QCD}}/(m_{A}c^{2})=1.132~(10)\times 10^{10}~{\rm MeV}~[8.8~{\rm per~mille}]$, with the value of the gravitational self-energy created by a pair of massless particles moving with this kinetic energy and separated by the distance, $\hbar c/\hat{E}$:
$f_{A}$ $=$ $G\frac{\left(\hat{E}/c^{2}\right)^{2}}{\hbar c/\hat{E}}$ (13)
$=$ $\frac{\hat{E}^{3}}{E_{\rm Pl}^{2}}.$
Solving Eq. (13), we find the predicted value for the kinetic energy at future infinity with per mille precision:
$\hat{E}({\rm CPT})$ $=$ $f_{A}({\rm CPT})^{1/3}E_{\rm Pl}^{2/3}$ (14)
$=$ $1.190~5~(35)\times 10^{18}~{\rm MeV}~[{\rm 2.9~per~mille}].$
Indeed, the value $\hat{E}({\rm CPT})\sim 10^{18}~{\rm MeV}$ is comparable to the “grand unification” energy scale, and, thus, supports the relation given in Eq. (12) between this kinetic energy and that of the Big Bang, $\check{E}$.
As a consistency check on the basic physical picture, we combine the relations given in Eq. (12) and Eq. (13) to derive a percent-level precise estimate for the value of the kinetic energy at the Bang:
$\check{E}({\rm CPT})$ $=$ $f_{A}({\rm CPT})^{5/3}E_{\rm Pl}^{-2/3}$ (15)
$=$ $107.6~(1.6)~{\rm MeV}~[{\rm 15~per~mille}].$
As expected, this value for the kinetic energy at the Big Bang in Eq. (15) is of the same order of magnitude as the estimate based on the value of the cosmological constant in Eq. (11). To go further, we use in Eq. (11) the actual values of the dimensionless numerically exact constants, $N_{U}=1/(2\pi^{2})=0.050~660~592$, and, $N_{n}=6\times 2!\zeta(3)/(\pi^{2}x_{k}^{3})=0.065~072~004$ (See Appendix A for the derivation of these values from the physical picture):
$\check{E}=104.218~7~(8)~{\rm MeV}~\left(\frac{\Omega_{\Lambda}h^{2}}{0.35}\right)^{1/6}~[7.3~{\rm ppm}],$ (16)
where the still rough but already pleasing agreement between the values in Eq. (15) and Eq. (16) suggests that our physical picture has brought us to the point where we can now estimate the cosmological parameters with confidence from first principles, and, in this way, resolve the Hubble tension.
3 0.5 eV QCD Axion Cosmological Parameters
Within the physical picture sketched in Section 2 (See also Appendix A), the most straightforward cosmological parameter to estimate is that of the 0.5 eV QCD axion dark matter, $\Omega_{A}h^{2}$, with the energy density, $u_{A}=n_{A}m_{A}c^{2}$, given by the axion rest-mass energy, $m_{A}c^{2}$, and by the axion number density, $n_{A}$:
$n_{A}$ $=$ $6\times n_{\gamma}$ (17)
$=$ $6\times\frac{2!\zeta(3)}{\pi^{2}}\left(\frac{kT_{\gamma}}{\hbar c}\right)^{3}$
$=$ $2~464.4~(1.6)~{\rm cm}^{-3}~[0.66~{\rm per~mille}],$
where the axion-to-photon number ratio, $n_{A}/n_{\gamma}=6$, is set to its value at the Big Bang within the physical picture sketched in Section 2 (For the derivation, see Appendix A). Combining this value for $n_{A}$ with the value of $m_{A}c^{2}$ from Eq. (6), we find a high-precision prediction from first principles for the cosmological parameter of the cold dark matter in the universe,
$\Omega_{A}h^{2}$ $=$ $\frac{8\pi G}{3c^{2}}\left(\frac{h}{H_{0}}\right)^{2}m_{A}c^{2}n_{A}$ (18)
$=$ $0.117~949~(78)~[0.66~{\rm per~mille}],$
where the exact value of the astronomical unit ${\rm AU}=149~597~870.700~{\rm km}$ (N. Capitaine et al., 2012) was used to derive the numerically exact value of the cosmological time, $h/H_{0}=({\rm AU}/100~{\rm km})(180/\pi)(60)^{2}\times 10^{6}~{\rm sec}=3.085~677~581\times 10^{17}~{\rm sec}.$
As noted in Section 1, the high-precision prediction from first principles for the cold dark matter cosmological parameter in Eq. (18) agrees to better than a percent with the best present estimate for this parameter within the standard $\Lambda{\rm CDM}$ cosmology using early-universe observations (Planck Collaboration et al., 2019, 2020)
$\Omega_{A}h^{2}(\Lambda{\rm CDM})=0.118~82~(86)~[7.2~{\rm per~mille}].$ (19)
Building on this apparent success of our physical picture for dark matter, dark energy, and the Big Bang in bridging the gap between early- and late-universe observations, a quick and dirty estimate of the cosmological constant from first principles is a natural next step. Combining the two estimates of the kinetic energy of radiation at the Big Bang, $\check{E}$, from Eq. (15) and Eq. (16), and using the high-precision first principles prediction for the value of the present expansion rate from Eq. (1), we get a percent-level precise estimate for the cosmological parameter, $\Omega_{\Lambda}$, for the cosmological constant
$\Omega_{\Lambda}({\rm CPT})$ $=$ $\left(0.35\right)\left(\frac{107.6~(1.6)}{104.2}\right)^{6}\frac{1}{(0.719~395)^{2}}$ (20)
$=$ $0.820~(73)~[90~{\rm per~mille}].$
Although necessarily rough, the estimate for the cosmological parameter of the cosmological constant in Eq. (20) still agrees to within twenty percent with the best present estimate for this parameter within the standard $\Lambda{\rm CDM}$ cosmology using early-universe observations (Planck Collaboration et al., 2019, 2020)
$\Omega(\Lambda{\rm CDM})=0.692~(5)~[7.2~{\rm per~mille}].$ (21)
To push the precision of the first principles estimate of the cosmological constant and to determine the last remaining cosmological parameter, $\Omega_{B}h^{2}$, that of the baryons, we must leverage the numerically exact value for the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}=0.399~286~(14)~[35~{\rm ppm}]$, given in Eq. (5), and derived in Sec. 4:
$\Omega_{B}h^{2}$ $=$ $\Omega_{A}h^{2}\times\left(\frac{\Omega_{B}h^{2}}{\Omega_{A}h^{2}}\right)$ (22)
$=$ $0.047~095~(31)~[0.66~{\rm per~mille}],$
where the value predicted within our physical picture in Eq. (18) was used for $\Omega_{A}h^{2}$. As expected, this high-precision first principles prediction for the baryon abundance in Eq. (22) differs sharply, by $190\sigma$, roughly a factor of two, from the best present estimate for this parameter within the standard $\Lambda{\rm CDM}$ cosmology using early-universe observations (Planck Collaboration et al., 2019, 2020)
$\Omega_{B}h^{2}(\Lambda{\rm CDM})=0.022~46~(13)~[5.8~{\rm per~mille}].$ (23)
As discussed in Section 1 (See Eq. (4)), this sharp discrepancy in the value for the baryon abundance predicted with our physical picture and the value predicted with the standard $\Lambda{\rm CDM}$ cosmology is the key to resolving the Hubble tension between early- and late-universe observations, since the values of the other cosmological parameters are basically the same in our physical picture as in $\Lambda{\rm CDM}$ cosmology.
To attain the promised 22 ppm precision for the value of $\Omega_{\Lambda}h^{2}$, it suffices to match the numerically exact baryon-to-axion energy density ratio, given in Eq. (5), with the value for this ratio computed using straightforward statistical mechanics in thermal equilibrium at the temperature of the Big Bang, $T_{\pi}=y_{T}(\chi_{\rm QCD})^{1/4}/k$, with the numerically exact dimensionless constant $y_{T}=(5/84)^{1/4}=0.493~938~274$ (See Appendix B for the derivation of $y_{T}$):
$\frac{\Omega_{B}h^{2}}{\Omega_{A}h^{2}}=\frac{1}{6}\left(\frac{m_{p}}{m_{A}}e^{-x_{p}}J(x_{p})+\frac{m_{n}}{m_{A}}e^{-x_{n}}J(x_{n})\right)\frac{1+\delta}{1+\epsilon},$ (24)
where the axion-to-photon number density ratio, $n_{A}/n_{\gamma}=6$, has been used (See Appendix A for derivation), and the nucleon mass ratio, $x_{N}=m_{N}c^{2}/(kT_{\pi})$, for N=p,nN=p,~n, determines the ratio of the trace of the baryon stress-energy tensor to the photon number density at the Big Bang, $x_{N}e^{-x_{N}}J(x_{N})$, in terms of the numerically exact integral,
$J(x_{N})=\frac{1}{2!\zeta(3)}\int_{0}^{\infty}dxe^{-x}x^{1/2}(x+2x_{N})^{1/2}.$ (25)
Using the value of the rest-mass of the 0.5 eV QCD axion from Eq. (6), derived in Appendix B, the matching of the baryon-to-axion energy density ratio in Eq. (24) with the numerically exact value predicted in Eq. (5) then determines the temperature of the Big Bang, $T_{\pi}$, to 0.17 ppm, and the topological susceptibility of the QCD vacuum, $\chi_{\rm QCD}$, to 0.68 ppm:
$kT_{\pi}$ $=$ $36.911~503~1~(63)~{\rm MeV}~[0.17~{\rm ppm}]$
$\sqrt{\chi_{\rm QCD}}$ $=$ $5~584.420~4~(19)~{\rm MeV^{2}}~[0.34~{\rm ppm}],$ (26)
where the value of $\sqrt{\chi_{\rm QCD}}$ agrees with the best present estimate from chiral perturbation theory (CPT) $\sqrt{\chi_{\rm QCD}}({\rm CPT})=5~690~(50)~{\rm MeV}^{2}~[8.8~{\rm per~mille}]$ (M. Gorghetto & G. Villadoro, 2019), to within two percent, but is nearly 26000 times more precise. At last, combining the Big Bang temperature from Eq. (3) with the 0.5 eV QCD axion rest-mass energy from Eq. (6) yields the cosmological parameter for the cosmological constant, $\Omega_{\Lambda}h^{2}$, to 22 ppm (See Appendix B for the derivation of the precision of the estimate and Appendix D for the derivation of the estimate itself):
$\Omega_{\Lambda}h^{2}$ $=$ $\frac{16\zeta(3)x_{k}^{6}y_{T}^{10}}{\pi^{3}}\frac{\left(m_{A}c^{2}\right)^{5}}{kT_{\pi}E_{\rm Pl}^{2}}\left(\frac{h}{H_{0}}\frac{1}{\hbar}\right)^{2}$ (27)
$=$ $0.352~460~0~(78)~[{\rm 22~ppm}],$
which combines with the estimates of $\Omega_{A}h^{2}$ in Eq. (18) and of $\Omega_{B}h^{2}$ in Eq. (22) to give the first-principles prediction for the present expansion rate of the universe in Eq. (1) using Eq. (4), and which agrees within two percent with the $\Lambda{\rm CDM}$ value in Eq. (21) for $\Omega_{\Lambda}$, given by
$\Omega_{\Lambda}$ $=$ $\frac{\Omega_{\Lambda}h^{2}}{h^{2}}$ (28)
$=$ $\frac{0.352~460~0}{(0.719~395~(76))^{2}}$
$=$ $0.681~04~(14)~[{\rm 0.21~per~mille}],$
where we used Eq. (27) for $\Omega_{\Lambda}h^{2}$ and Eq. (1) for $h$.
4 0.5 eV QCD Axion-to-Baryon Energy Density Ratio
To derive the numerically exact value of the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}$, given in Eq. (5), which is the key to the proposed resolution of the Hubble Tension, presented in Sections 1, 2, and 3, we use the cyclic cosmology matching condition (See Appendix C):
$\left(\frac{T_{\pi}}{T_{\gamma}}\right)^{3}\left(\frac{\check{E}}{f_{A}}\right)^{2}\left(\frac{1+\epsilon}{1+\delta}\right)\Omega_{B}h^{2}=4\left(\frac{\hat{E}}{f_{A}}\right)^{2}\Omega_{\Lambda}h^{2},$ (29)
which gives an independent relationship between the baryon energy density and the cosmological constant. To eliminate the cosmological constant and bring in the axion energy density, we use the expression for the cosmological constant, $\Lambda=8\pi Gu_{\Lambda}/c^{4}$, in terms of dark energy density, $u_{\Lambda}=U_{\Lambda}n_{\Lambda}$, with the dark energy of particle-antiparticle pairs at the Big Bang, $U_{\Lambda}$, given in Eq. (8) with $\check{E}$ for the kinetic energy, and with the particle-antiparticle pair density, $n_{\Lambda}=(T_{\pi}/T_{\gamma})^{3}n_{A}$, set to the axion number density at the Big Bang (See Appendix A),
$\Omega_{\Lambda}h^{2}$ $=$ $\frac{8\pi G}{3c^{2}}\left(\frac{h}{H_{0}}\right)^{2}\left(N_{U}G\frac{(\check{E}/c^{2})}{\hbar c/\check{E}}\right)\left(\frac{T_{\pi}}{T_{\gamma}}\right)^{3}n_{A}$ (30)
$=$ $\left(\frac{1}{2\pi^{2}}\frac{\check{E}^{3}}{E_{\rm Pl}^{2}}\right)\left(\frac{T_{\pi}}{T_{\gamma}}\right)^{3}\left(\frac{\Omega_{A}h^{2}}{m_{A}c^{2}}\right),$
where the axion energy density in the present universe, $u_{A}=m_{A}c^{2}n_{A}$, was used in the last line, and the value of the dimensionless constant, $N_{U}=1/(2\pi^{2})$, is derived in Appendix A. Combining the cyclic cosmology matching condition in Eq. (29) with the expression for the cosmological constant in terms of the dark energy density at the Big Bang in Eq. (30), we derive the remarkable, numerically exact baryon-to-axion ratio for energy density in the present universe given in Eq. (5)
$\frac{\Omega_{B}h^{2}}{\Omega_{A}h^{2}}$ $=$ $\left(\frac{\left(T_{\pi}/T_{\gamma}\right)^{3}\Omega_{B}h^{2}}{\Omega_{\Lambda}h^{2}}\right)\left(\frac{\Omega_{\Lambda}h^{2}}{\left(T_{\pi}/T_{\gamma}\right)^{3}\Omega_{A}h^{2}}\right)$ (31)
$=$ $\frac{2}{\pi^{2}}\left(\frac{\check{E}}{m_{A}c^{2}}\right)\left(\frac{\hat{E}^{2}}{E_{\rm Pl}^{2}}\right)\left(\frac{1+\delta}{1+\epsilon}\right)$
$=$ $\frac{2}{\pi^{2}}\left(\frac{\check{E}^{2}}{m_{A}c^{2}f_{A}}\frac{f_{A}}{\check{E}}\right)\left(\frac{f_{A}}{\hat{E}}\right)\left(\frac{1+\delta}{1+\epsilon}\right)$
$=$ $\frac{2}{\pi^{2}}\left(\frac{\check{E}}{\sqrt{\chi_{\rm QCD}}}\right)^{2}\left(\frac{f_{A}^{2}}{\check{E}\hat{E}}\right)\left(\frac{1+\delta}{1+\epsilon}\right)$
$=$ $\frac{2}{\pi^{2}}\left(x_{k}y_{T}\right)^{2}\left(\frac{1+\delta}{1+\epsilon}\right),$
where we have used the see-saw relation, $\check{E}\hat{E}=f_{A}^{2}$, from Eq. (12) in the last line and the dark energy condition at future infinity, $f_{A}=\hat{E}^{3}/E_{\rm Pl}^{2}$, from Eq. (13) in the third line.
5 0.5 eV QCD Axion Structure Formation
Figure 1: Cooling of 0.5 eV QCD axion dark matter in the early universe shows a cross-over in the adiabatic compressibility, $dp_{A}/du_{A}$ (purple crosses), from the radiation-like value, $dp_{A}/du_{A}\approx 1/3$, at large redshift, to the matter-like value, $dp_{A}/du_{A}\approx 0$, at small redshift, starting right around the redshift of “matter-radiation equality,” $z_{\rm eq}=3~376~(19)$ (vertical orange arrow), predicted by the standard cold dark matter cosmology with cosmological constant ($\Lambda{\rm CDM}$) using observations of the early universe (Planck Collaboration et al., 2019, 2020).
The formation of large-scale astrophysical structures in the early universe is a key test of any proposal for the nature of dark matter, dark energy, and the Big Bang, including 0.5 eV QCD axion cyclic cosmology. Historically, the now-standard cold dark matter cosmology first emerged in the context of concerns about the relative lack of large-scale anisotropy in the cosmic microwave background (CMB), compared to expectations based on the then-standard baryon-based Big Bang cosmology (P. J. E. Peebles, 2020, pg. 301–302). In fact, Peebles first proposed a hot dark matter cosmology (P. J. E. Peebles, 1982a), before turning to cold dark matter (P. J. E. Peebles, 1982b), with the shared goal, in both cases, of “washing out” large-scale CMB anisotropy and, in this way, resolving the tension between the then-standard cosmology and early-universe observations by D. J. Fixsen et al. (1983).
Figure 2 shows the cross-over in the equation of state that relates the pressure, $p_{A}(z)$, and the energy density, $u_{A}(z)$, for 0.5 eV QCD axion dark matter in thermal equilibrium with the cosmic background radiation at temperature, $(1+z)T_{\gamma}$ (P. J. E. Peebles, 1993, pgs. 134–135), where $z$ is the redshift and $T_{\gamma}$ is the temperature of the CMB. At large redshift, $z\gg 10^{4}$, the adiabatic axion compressibility, $dp_{A}/du_{A}\approx 1/3$, takes the radiation-like value, reflecting a light-like equation of state, $p_{A}(z)=u_{A}(z)/3$ (P. J. E. Peebles, 1993, pg. 139), while at small redshift, $z\ll 10^{2}$, the value crosses over to matter-like, $dp_{A}/du_{A}\approx 0$. Crucially, the cross-over begins at roughly the best present estimate of the redshift for “matter-radiation equality,” $z_{\rm eq}=3~376~(19)$ (Planck Collaboration et al., 2019, 2020), within the standard cold dark matter cosmology.
As the 0.5 eV QCD axion dark matter equation of state crosses over to matter-like at small redshifts from radiation-like at large redshifts, starting around $z_{\rm eq}$, as shown in Fig. 2, the dynamics of the expansion change to “matter-dominated” from “radiation-dominated.” Indeed, with six axions for each photon made in the cosmic background radiation (See Section 3 and Appendix A), axions dominate the expansion dynamics on both sides of “matter-radiation equality.” The cross-over in the axion equation of state (Fig. 2), beginning around $z_{\rm eq}$, drives the cross-over in the character of the expansion which then drives the large-scale structure formation seen by early-universe observations, just as the change from “radiation-domination” to “matter-domination” drives such structure to form in the standard cold dark matter cosmology (P. J. E. Peebles, 1993, pgs. 624–625).
The rough agreement, shown in Fig. 2, between the best present observational value of the redshift where large-scale structure formation begins, $z_{\rm eq}$, and the predicted value for this redshift based on the cross-over in the equation of state for the 0.5 eV QCD axion dark matter, is a crucial test for the proposed cosmology. Together with the striking agreement between the predicted and observed values for the present energy density of cold dark matter in the universe (See Section 3), this agreement on large-scale structure formation provides compelling evidence that the proposed 0.5 eV QCD axion dark matter cosmology is compatible with early-universe observations. In this way, we deepen the proposed resolution of the Hubble tension, sketched in Section 1, by ensuring that the agreement between the standard $\Lambda{\rm CDM}$ cosmology and early-universe observations is not spoiled by the switch to “cooling” dark matter, in the form of 0.5 eV QCD axion dark matter with the equation of state shown in Fig. 2.
6 0.5 eV QCD Axion Nucleosynthesis
Figure 2: Cooling of the universe in the first three minutes after the Big Bang for a mix with six 0.5 eV QCD axions for each photon (curved line) and for the standard mix of photons and light leptons (filled and unfilled boxes) both reach the freeze-out of the neutron-to-proton ratio (horizontal line with dots) around the same time, but the standard mix (P. J. E. Peebles, 1966; G. Steigman, 2007) takes roughly twice as long to start forming helium nuclei (horizontal line with boxes).
As the expanding universe cools in the first three minutes after the Big Bang (Fig. 3), the mix of particles with a light-like equation of state sets the time, $t_{d}$, that it takes to “turn off” the photodisintegration of the deuteron, $d$, so that the nuclei of light chemical elements beyond hydrogen, mainly helium, can form (P. J. E. Peebles, 1993, pgs. 184–186). Remarkably, the resulting “primordial” helium abundance (by mass), $Y_{\rm P}\approx 0.3$, does not depend much on the baryon number density, $n_{B}(T_{d})$, at the “turn-off” temperature, $T_{d}\approx 10^{9}~{\rm K}$ (P. J. E. Peebles, 2020, pg. 129). Instead, the main constraints on baryon abundance from Big Bang nucleosynthesis (BBN) come from comparisons with observations of the “primordial” deuterium abundance (by number), $n_{d}/n_{p}=2.55~(3)\times 10^{-5}$ (R. J. Cooke et al., 2018), based on quasar absorption line spectroscopy.
Indeed, doubling the baryon number density, $n_{B}(T_{d})$, compared to the standard cold dark matter cosmology value, as we propose to do to resolve the Hubble tension (See Section 1 and 3), ensures a nearly complete conversion of deuterons into helium, and, thus, a nearly complete absence of primordial deuterium (R. J. Cooke et al., 2018), for the timing of the turn-off, $t_{d}(\rm{\Lambda CDM})\approx 200~{\rm sec}$ (P. J. E. Peebles, 1966; G. Steigman, 2007), set by the standard mix of light leptons and photons. Yet, with the appropriate mix of six axions for every photon made in the Big Bang (See Section 3 and Appendix A), the turn-off time, $t_{d}\approx 100~{\rm sec}$, comes nearly twice as fast as it does with the standard mix (See Fig. 3). Thus, the value of the product that determines the deuterium abundance, $n_{B}(T_{d})\langle\sigma v\rangle t_{d}\sim 1$, lands in about the same place, preserving the agreement with observations, where $\langle\sigma v\rangle$ is the thermal average of the product of the neutron capture cross-section, $\sigma$, with the nucleon relative velocity, $v$, (P. J. E. Peebles, 1993, pg. 141–142).
Still, the change in the time of the turn-off, $t_{d}$, has observable consequences due to the change in the gap in time, $t_{d}-t_{n}$, between the start of BBN, at $t_{d}$, and the “freeze-out,” at $t_{n}$, of the weak nuclear reactions between light leptons and nucleons (Fig. 3). During this gap, the decay of free neutrons with life-time, $\tau_{n}\approx 878~{\rm sec}$ (P. J. Mohr et al., 2025), cuts down the neutron-to-proton ratio (P. J. E. Peebles, 1993, pg. 184–186). By cutting $t_{d}$, while leaving $t_{n}$ more or less unchanged, we shorten the gap, $t_{d}-t_{n}$, by about ten percent of the free neutron lifetime, $\tau_{n}$, so that about ten per cent more neutrons make it to the start of BBN, at $t_{d}$, after which essentially all the surviving free neutrons end up as primordial helium (P. J. E. Peebles, 2020, pg. 128–129): We thus expect a roughly ten percent larger value for the primoridal helium abundance, $Y_{\rm P}\approx 0.27$, compared to the best present estimate with the standard mix of light leptons and photons, $Y_{\rm P}({\rm\Lambda CDM})\approx 0.25$ (C. Pitrou et al., 2018; S. Gariazzo et al., 2022).
7 0.5 eV QCD Axion Initial Fluctuations
More than five decades ago, E. R. Harrison (1970) and Y. B. Zeldovich (1972) argued for a scale-invariant initial spectrum of fluctuations in the energy density of the radiation created in the Big Bang, $P(k)=Ak^{n_{s}}$, where the spectral index value, $n_{s}=1$, ensures scale-invariance, and the constant, $A$, is independent of the co-moving wavevector of the fluctuation, $k$. Within the standard $\Lambda{\rm CDM}$ cosmology, the best present estimate from early-universe observations for the value of the spectral index, $n_{s}({\rm\Lambda CDM})=0.9677~(37)$ (Planck Collaboration et al., 2019, 2020), falls $8.8\sigma$ below the scale-invariant value, $n_{s}=1$. By contrast, the conformal symmetry of the cross-over from future infinity to the next big bang in 0.5 eV QCD axion cyclic cosmology (See Appendix C) requires the scale-invariant value, $n_{s}=1$, in agreement with the argument of E. R. Harrison (1970) and Y. B. Zeldovich (1972), but in sharp tension ($8.8\sigma$) with the standard $\Lambda{\rm CDM}$ value.
8 0.5 eV QCD Axion Astrophysical Constraints
Figure 3: Axion-photon coupling grows with initial core helium mass fraction in simulations of stellar evolution for horizontal branch stars constrained by observations of globular clusters (curved line with dots) (A. Ayala et al., 2014). The predicted value for the axion-photon coupling strength (horizontal lines with boxes), given in Eq. (7), combines with this empirical-computational relation to give a prediction for the helium mass fraction (box with whiskers) that differs by $5.0\sigma$ from the primordial helium abundance expected within the standard cold dark matter cosmology with cosmological constant ($\rm\Lambda CDM$}) based on simulations of Big Bang nucleosynthesis (thick vertical line) (S. Gariazzo et al., 2022; C. Pitrou et al., 2018), but that agrees rather well with the best present estimates for the initial core helium mass fraction in the Sun (pair of thin vertical lines) (L. Piersanti et al., 2007; A. M. Serenelli & S. Basu, 2010).
Attempts to “constrain” or “exclude” 0.5 eV QCD axion dark matter using astrophysical phenomena can be grouped by the particle to which the axion couples to create the effect: photon, electron, or nucleon. For axion-photon coupling, the most recent attempts are based on the spontaneous two-photon decay of 0.5 eV QCD axions in the dark matter “halo” of the Milky Way (E. Pinetti, 2025; R. Janish & E. Pinetti, 2025; S. Roy et al., 2025), while a more established approach looks for axion production from gamma rays in the cores of horizontal branch stars within globular clusters (A. Ayala et al., 2014). Efforts involving axion-electron coupling go back five decades (K. Sato & H. Sato, 1975; K. Sato, 1978), focused mainly on red giants and white dwarfs, while attempts invoking axion-nucleon coupling have instead looked at pulsars and supernova 1987A (See A. Caputo & G. Raffelt (2024) for a recent review).
Figure 4 shows the empirical-computational relation between the axion-photon coupling strength, $g_{A\gamma\gamma}=g_{10}\times 10^{-10}~{\rm GeV}^{-1}$, and the initial core helium mass fraction, $Y_{\rm HB}$, for stars that are now on the horizontal branch (HB) of the color-magnitude diagram for globular clusters (A. Ayala et al., 2014):
$R=6.26Y_{\rm HB}-0.41g_{10}^{2}-0.12,$ (32)
where $R=N_{\rm HB}/N_{\rm RGB}=1.39~(3)$ was the observed ratio of the number of HB stars, $N_{\rm HB}$, to the number of stars on the red giant branch (RGB) of the color-magnitude diagram, $N_{\rm RGB}$, in 39 globular clusters. Combining the empirical-computational relation in Eq. (32) with the predicted value for the axion-photon coupling in Eq. (7), we find a percent-level precise estimate for the initial core helium mass fraction in these HB stars that formed in the first billion years after the Big Bang:
$Y_{\rm HB}=0.272~(5)~[18~{\rm per~mille}],$ (33)
which differs by $5\sigma$ with the best present estimate of the “primordial” helium abundance, $Y_{P}({\rm\Lambda CDM})=0.247~09~(17)~[0.69~{\rm per~mille}]$ (C. Pitrou et al., 2018; S. Gariazzo et al., 2022), within the standard cold dark matter cosmology, but that roughly matches the best present estimates for the initial core helium mass fraction in the Sun, $Y_{\odot}=0.273~(6)~[{\rm 22~per~mille}]$ (A. M. Serenelli & S. Basu, 2010; L. Piersanti et al., 2007). As discussed in Section 6, the mix of axions and photons that dominates the expansion of the universe leading up to Big Bang nucleosynthesis, within the proposed physical picture, yields about ten percent more primordial helium, $Y_{P}\approx 0.27$, than the standard mix of light leptons and photons, and, in this way, resolves the $5\sigma$ tension between $Y_{\rm HB}$ in Eq. (33) and $Y_{P}({\rm\Lambda CDM})$.
Similarly, the recent attempts to “rule out” axion dark matter with the rest-mass energy in Eq. (6) and the axion-photon coupling in Eq. (7) using “blank-sky” observations of the dark matter “halo” of the Milky Way at the near-infrared wavelength $2\lambda_{A}=2h/m_{A}c\approx 4.92~{\rm\mu m}$, corresponding to the decay of the axion into a pair of photons, subtract a “smooth astrophysical background” by fitting a quartic polynomial to the near-infrared spectrum within a window of width $\Delta\lambda\approx 0.2~{\rm\mu m}$ (E. Pinetti, 2025). However, the proposed physical picture implies that 0.5 eV QCD axion dark matter is in thermal equilibrium with the cosmic microwave background (See Section 3 and Appendix A), and, thus, the axions in the dark matter “halo” of the Milky Way have velocity dispersion, $\sigma_{v}$, fixed by the CMB temperature, $T_{\gamma}$, and by the 0.5 eV QCD axion dark matter rest-mass energy, $m_{A}c^{2}$, leading to an expected line-width, $2\Delta\lambda_{A}$, that is roughly the same size as the window used to subtract the “smooth astrophysical background:”
$2\Delta\lambda_{A}$ $=$ $2\lambda_{A}\frac{\sigma_{v}}{c}$ (34)
$=$ $2\lambda_{A}\sqrt{\frac{3kT_{\gamma}}{m_{A}c^{2}}}$
$\approx$ $0.2~{\rm\mu m}.$
Instead of “constraining” the 0.5 eV QCD axion, the analysis by E. Pinetti (2025) in effect rules out the standard cold dark matter velocity dispersion, $\sigma_{v}({\rm\Lambda CDM})\approx 200~{\rm km~sec^{-1}}$ (M. W. Goodman & E. Witten, 1985), which is based on the assumption of gravitational equilibrium with “bright” matter, such as OB associations, H II regions, and 21 cm clouds, and that is much smaller than the velocity dispersion, $\sigma_{v}\approx 11,000~{\rm km~sec^{-1}}$, given in Eq. (34) by thermal equilibrium with the CMB, within the proposed physical picture.
Turning to axion-electron coupling, the key point is that there is no “tree-level” or “bare” coupling between the axion and electron (R. D. Peccei, 1989), just as there is no such coupling between the neutral pion, $\pi^{0}$, and the electron (T. Husek, 2024). Instead, the “leading order” axion-electron coupling appears only at “loop-level,” and is suppressed by the small factor, $(\frac{\alpha}{2\pi})^{2}\sim 10^{-6}$ (P. J. Mohr et al., 2025), compared to the “bare” value used to derive astrophysical “constraints” and “exclusions” (K. Sato & H. Sato, 1975; K. Sato, 1978; A. Caputo & G. Raffelt, 2024). The upshot is that the effect of axions on the astrophysical phenomena, such as white dwarf cooling and red giant luminosity, is smaller than the astrophysical uncertainties entering the analysis, and, thus, no significant “constraints” or “exclusions” emerge from the comparison with astrophysical observations.
Finally, regarding the axion-nucleon coupling, a dense bunch of nucleons makes a lot of 0.5 eV QCD axions due to this coupling, but, by the same token, such a dense nuclear soup is “axionically thick,” in the sense that a 0.5 eV QCD axion made by one nucleon is unlikely to “escape” the concentration without coupling to another nearby nucleon, as pointed out by M. S. Turner (1988), in the analysis of the effect of axions on supernova 1987A. Remarkably, there is a relatively narrow window of QCD axion rest-mass energy, $0.05~{\rm eV}\leq m_{A}c^{2}\leq 5~{\rm eV}$ (W. Yin, 2023, pg. 42), that brackets the predicted value in Eq. (6), in which the effects of the QCD axions completely dominate the flow of energy within the dense nuclear soup, in sharp contrast with the traditional assumption that axions play essentially no role in this flow, and this contrast is often expressed as an “exclusion” of the 0.5 eV QCD axion. However, given the apparent success of 0.5 eV QCD axion dark matter in resolving the Hubble tension, which is the main focus of this Letter, one is instead led, on the basis of this analysis of the axion-nucleon coupling, to question the traditional assumptions regarding the flow of energy within compact astrophysical bodies, such as pulsars, as well as the flow within the events that are thought to form them, such as Type II supernovae.
9 Conclusions
Among astrophysicists during the past forty years, three main “pillars” have come to be viewed as supporting the now-standard cold dark matter cosmology ($\rm\Lambda CDM$}): the anisotropy of the cosmic microwave background (CMB), the “primordial” abundance of chemical elements with low atomic mass whose nuclei were synthesized in the Big Bang (BBN), and the redshift-magnitude relation of Type Ia supernovae (P. J. E. Peebles, 2020, pg. 323–340). The physical picture of dark matter, dark energy, and the Big Bang, presented in this Letter, rests on the same “pillars” of support, but does so more naturally. In particular, the proposed physical picture naturally resolves the now more than $5\sigma$ tension in the present expansion rate of the universe between $\rm\Lambda CDM$} and “late” universe observations, and does so in favor of the observations.
Still, most of the cosmological parameters come out with around the same values as they do in $\rm\Lambda CDM$}. The energy density of cold dark matter in the present universe and the cosmological constant, for example, land within the uncertainties of the best present $\rm\Lambda CDM$} values from fits to early universe observations. However, it is crucial to emphasize that these essentially shared values arise as first-principles predictions, not fits to observations, within the proposed physical picture.
Similarly, the redshift of “matter-radiation equality” and the “primordial” deuterium abundance come out about the same as $\rm\Lambda CDM$}. This redshift is the critical “epoch” in the early universe where large-scale structure begins to form, and its value determines the location of the main “peak” in the angular power spectrum of the CMB fluctuations. Likewise, the primordial deuterium abundance is a key indicator of the timing and the conditions in the Big Bang when “light” element nucleosynthesis begins.
In conclusion, what the standard $\rm\Lambda CDM$} cosmology does well, the present physical picture does just as well or better. More sharply, the present physical picture provides an actual physical picture for the main ingredients of $\rm\Lambda CDM$}: dark matter, dark energy, and the Big Bang. In this sense, $\rm\Lambda CDM$} can be viewed, as it was intended to be viewed (P. J. E. Peebles, 1993, pg. 621–622), as a way of stimulating “the development of ideas on how to find and test models for the origin of [large-scale] structure,” rather than as an actual physical picture for the nature of dark matter (or anything else).
Moving forward, the sharp difference in the baryon abundance, roughly $190\sigma$, gives about a factor of two more baryons with the present physical picture. At the same time, the mix of six axions for every photon made in the Big Bang, within this picture, speeds up the cooling off of the universe in the first three minutes. The combination of these two sharp differences leads to about the same deuterium abundance but about $5\sigma$ more helium compared to $\rm\Lambda CDM$}, we have argued, and these loose arguments, in part based on astrophysical observations of globular clusters, must be confirmed by actual BBN calculations within the proposed physical picture.
In a similar vein, an actual calculation of the spectrum of CMB fluctuations within the proposed physical picture must be done. We have argued, indirectly, that the location of the main peak in the spectrum should come out about the same as it does in $\rm\Lambda CDM$} due to the coincidence of the redshift where the 0.5 eV QCD axion dark matter equation of state crosses over from radiation-like to matter-like, on the one hand, and, on the other, the redshift of “matter-radiation equality” within $\rm\Lambda CDM$}. Along similar lines, there is reason to believe that the sharp difference in baryon abundance, roughly $190\sigma$, combines with the sharp difference in the “primordial” spectral index, nearly $9\sigma$, to give basically the same spectrum as the best fit within $\rm\Lambda CDM$}, so that the agreement of this best fit with early universe observations is retained in the present physical picture, albeit with first-principles predictions taking the place of the key fit parameters.
Finally, the straightforward resolution of the Hubble tension using 0.5 eV QCD axion cyclic cosmology, presented in this Letter, suggests that we must “re-visit” many of the conventional assumptions about the way dark matter interacts with light energy and “bright” matter. For much of the past half century, it has been conventional to assume that dark matter “tracks” bright matter in gravitational equilibrium within a “halo” around the Milky Way galaxy, but the lack of a sharp two-photon decay “line” from 0.5 eV QCD axion dark matter in the infrared spectrum of this halo, suggests that dark matter is, instead, in thermal equilibrium with the photons of the CMB, as predicted by the proposed physical picture. For an equally long span of time, axions have been assumed to play essentially no role in the flow of energy within compact astrophysical bodies, such as pulsars, and within Type II supernovae, such as supernova 1987A, and, thus, it may be time to re-visit those assumptions, in light of the resolution of the Hubble tension presented in this Letter.
Daria Mazura provided critical and constant support without which the work reported here would not have been possible.
Appendix A 0.5 eV QCD Axion Self-Energy and Number Density
Within the physical picture presented in Section 2, the gravitational self-energy of 0.5 eV QCD axion dark matter, $U_{\Lambda}=N_{U}G(E/c^{2})^{2}/R$, given in Eq. (8), is generated by the kinetic energy, $E$, of a massless particle and anti-particle, each with de Broglie wavelength, $\lambda=hc/E=2\pi R$. The value of the dimensionless constant, $N_{U}=1/(2\pi^{2})$, then follows from two assumptions. First, the axion is “spin-less,” in the sense that the plane of motion of the particle and anti-particle has equal probability amplitude to be oriented in any direction in three-dimensional space, and, second, the three-dimensional vector, $\bf r$}, separating the particle and anti-particle, is related to twice the de Broglie wavelength, $D=2\lambda$, by the familiar geometric relation, $1/D=\langle 1/|{\bf r}|\rangle_{\Omega}$, for the angle-averaged plane-projected (“line-of-sight”) distance (V. Barger & M. Olsson, 1995, pgs. 338–339):
$\frac{1}{D}$ $=$ $\left\langle\frac{1}{|\bf{r}|}\right\rangle_{\Omega}$ (A1)
$=$ $\frac{1}{4\pi}\int\frac{d\Omega}{r\sin\theta}$
$=$ $\frac{1}{4\pi r}\int_{0}^{\pi}d\theta\int_{0}^{2\pi}d\phi$
$=$ $\frac{\pi}{2r},$
where, $r$ is the three-dimensional separation, $\Omega$ is the solid angle on the two-dimensional sphere of orientations of the orbital plane of the particle-antiparticle pair relative to some fixed “line-of-sight,” $\theta$ is the polar angle (co-altitude) of the line-of-sight to the normal vector of the plane, and $\phi$ is the azimuthal angle of this normal vector with respect to the line-of-sight.
Using the geometric identity in Eq. (A1), it is then straightforward to derive the gravitational self-energy of the 0.5 eV QCD axion dark matter:
$U_{\Lambda}$ $=$ $G\frac{(E/c^{2})^{2}}{r}$ (A2)
$=$ $\frac{2}{\pi}G\frac{(E/c^{2})^{2}}{D}$
$=$ $\frac{1}{2\pi^{2}}G\frac{(E/c^{2})^{2}}{R}.$
Comparing the last line of Eq. (A2) with the expression for the gravitational self-energy in Eq. (8), we find the value of the dimensionless constant, $N_{U}=1/(2\pi^{2})$. We emphasize that this result requires that the 0.5 eV QCD axion dark matter be a spin-less bound-state of particle and anti-particle, and, thus, the axion must obey Bose-Einstein statistics (L. D. Landau & E. M. Lifshitz, 1958, pg. 205), just as photons do (P. J. E. Peebles, 1993, pg. 134).
At the Big Bang, the average kinetic energy, $\check{E}=x_{k}kT_{\pi}\approx 104~{\rm MeV}$, given in Eq. (16), is much larger than the 0.5 eV QCD axion dark matter rest-mass energy, $m_{A}c^{2}=0.5~{\rm eV}=5\times 10^{-7}~{\rm MeV}$, given in Eq. (6). The axion number density at the Big Bang, $n_{A}(T_{\pi})$, is then determined by the Planck distribution function, just as the photon number density at the Big Bang, $n_{\gamma}(T_{\pi})$, is so determined (P. J. E. Peebles, 1993, pg. 134–137, 158–159), with the result
$n_{A}(T_{\pi})$ $=$ $\frac{g_{A}}{2}\frac{2!\zeta(3)}{\pi^{2}}\left(\frac{kT_{\pi}}{\hbar c}\right)^{3}$ (A3)
$=$ $6n_{\gamma}(T_{\pi})$
$=$ $\left(\frac{T_{\pi}}{T_{\gamma}}\right)^{3}n_{A},$
where $n_{A}$ is the axion number density in the present universe, $T_{\gamma}$ is the temperature of the cosmic microwave background, and the statistical degeneracy factor, $g_{A}=12$, counts the twelve “flavors” of quarks and leptons in the standard model of particle physics (L. M. Lederman & D. N. Schramm, 1989, pg. 128). Each flavor, $M$, of quark and lepton binds with an anti-particle of the same flavor and helicity to make the axion, $A_{M}$, with this flavor and no spin, such that the axion number density at the Big Bang is given by the expression in Eq. (9), with the value of the dimensionless constant, $N_{n}=6\times 2!\zeta(3)/(\pi^{2}x_{k}^{3})$, fixed by the result given above in Eq. (A3).
Appendix B 0.5 eV QCD Axion Rest-Mass and Big Bang Temperature
The predicted value for the rest-mass energy, $m_{A}c^{2}$, of the 0.5 eV QCD axion dark matter, given in Eq. (6), follows from the analysis of the baryon-to-axion energy density ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}$, presented at the end of Section 3, once a relation can be found with the temperature of the Big Bang, $T_{\pi}$. Both of the quantities, $m_{A}c^{2}$ and $kT_{\pi}$, enter the statistical mechanical expression for this ratio, given in Eq. (24), through their ratios with the nucleon rest-mass energies, $m_{N}c^{2}$, for N=n,pN=n,~p. Given an independent relation between $m_{A}c^{2}$ and $kT_{\pi}$, then, the matching of this expression to the numerically exact value for the ratio, $\Omega_{B}h^{2}/\Omega_{A}h^{2}$, given in Eq. (5) and derived in Section 4, determines both quantities with a precision that is limited only by that of this independent relation.
The starting point for finding such an independent relation between $m_{A}c^{2}$ and $kT_{\pi}$ is the relation between the axion decay constant, $f_{A}$, and the kinetic energy, $\hat{E}$, of the massless particle and anti-particle within the axion at future infinity, given in Eq. (13), combined with the “see-saw” relation for the value of this kinetic energy, $\check{E}=x_{k}kT_{\pi}$, at the Big Bang, given in Eq. (12):
$f_{A}$ $=$ $\left(\frac{f_{A}^{2}}{x_{k}kT_{\pi}}\right)^{3}\left(E_{\rm Pl}\right)^{-2}$ (B1)
$=$ $f_{A}\left(\frac{\sqrt{\chi_{\rm QCD}}}{m_{A}c^{2}}\right)^{5}\left(x_{k}kT_{\pi}\right)^{-3}\left(E_{\rm Pl}\right)^{-2}$
$=$ $f_{A}\frac{1}{x_{k}^{3}y_{T}^{10}}\left(kT_{\pi}\right)^{7}\left(m_{A}c^{2}\right)^{-5}\left(E_{\rm Pl}\right)^{-2}$
where the topological susceptibility of the QCD vacuum, $\chi_{\rm QCD}=\left(m_{A}c^{2}f_{A}\right)^{2}$, gives the temperature of the Big Bang, $kT_{\pi}=y_{T}(\chi_{\rm QCD})^{1/4}$, with the dimensionless constant, $y_{T}=(5/84)^{1/4}$, to be derived at the end of this appendix. Solving for $m_{A}c^{2}$ in Eq. (B1), we find the desired independent relation with $kT_{\pi}$:
$m_{A}c^{2}=x_{k}^{-3/5}y_{T}^{-2}\left(kT_{\pi}\right)^{7/5}\left(E_{\rm Pl}\right)^{-2/5}.$ (B2)
The precision of this relation is set by the 22 ppm uncertainty in the best present value for the universal gravitational constant, $G=6.674~08~(15)\times 10^{-8}~{\rm g^{-1}~cm^{3}~s^{-2}}$ (P. J. Mohr et al., 2025), and, we obtain the precision of the determinations of $m_{A}c^{2}$ and $kT_{\pi}$ by simply varying the value of $E_{\rm Pl}$ in Eq. (B2), within the 11 ppm window set by $\Delta G/G$, and reading off the changes, $\Delta m_{A}c^{2}$ and $\Delta kT_{\pi}$, with the following empirical-computational results:
$\frac{\Delta m_{A}c^{2}}{m_{A}c^{2}}$ $=$ $24\times 10^{-3}~\frac{\Delta G}{G}={\rm 0.53~ppm}$
$\frac{\Delta kT_{\pi}}{kT_{\pi}}$ $=$ $7.7\times 10^{-3}~\frac{\Delta G}{G}={\rm 0.17~ppm}$
$\frac{\Delta\chi_{\rm QCD}}{\chi_{\rm QCD}}$ $=$ $4\frac{\Delta kT_{\pi}}{kT_{\pi}}={\rm 0.68~ppm}$
$\frac{\Delta f_{A}}{f_{A}}$ $=$ $\left|2\frac{\Delta kT_{\pi}}{kT_{\pi}}-\frac{\Delta m_{A}c^{2}}{m_{A}c^{2}}\right|={\rm 0.19~ppm}$
$\frac{\Delta\Omega_{\Lambda}h^{2}}{\Omega_{\Lambda}h^{2}}$ $=$ $\frac{\Delta G}{G}={\rm 22~ppm},$ (B3)
where the precision of the derived quantities, $\chi_{\rm QCD}=y_{T}^{-4}(kT_{\pi})^{4}$, $f_{A}=\sqrt{\chi_{\rm QCD}}/(m_{A}c^{2})=y_{T}^{-2}(kT_{\pi})^{2}(m_{A}c^{2})^{-1}$, and, $\Omega_{\Lambda}h^{2}\propto G$, follows from straight-forward error-propagation (P. J. Mohr et al., 2025).
Finally, we derive the relation, $kT_{\pi}=y_{T}\left(\chi_{\rm QCD}\right)^{1/4}$, with $y_{T}=(5/84)^{1/4}$, between the Big Bang temperature, $T_{\pi}$, and the QCD topological susceptibility, $\chi_{\rm QCD}$. The starting point is simply the identification, on the one hand, of the kinetic energy density of the photons and axions created in the Big Bang, $(6+1)(T_{\pi}/T_{\gamma})^{4}(\hbar c)^{3}u_{\gamma}$, where $(\hbar c)^{3}u_{\gamma}=(\pi^{2}/15)(kT_{\gamma})^{4}$ (P. J. E. Peebles, 1993, pg. 137) gives the energy density of the cosmic microwave background in the present universe (See Appendix A for the statistical degeneracy factor, $g_{A}/2=6$, that gives the axion-to-photon energy density ratio at the Big Bang), and, on the other hand, of the potential energy density, $\chi_{\rm QCD}(\Delta\theta_{A}/6)^{2}$ (G. ’t Hooft, 1976a, b), released by the QCD vacuum when the axion “angle” jumps by the amount, $\Delta\theta_{A}=\pi$, from the value, $\theta_{A}=\pi$, before the Big Bang to the value, $\theta_{A}=0$, after the Big Bang (R. D. Peccei & H. R. Quinn, 1977a, b):
$7\frac{\pi^{2}}{15}\left(kT_{\pi}\right)^{4}=\chi_{\rm QCD}\left(\frac{\pi}{6}\right)^{2}.$ (B4)
Solving Eq. (B4) for $kT_{\pi}$, we find the desired relation.
Appendix C 0.5 eV QCD Axion Cyclic Cosmology
[…]
$\hat{\ell}$ $=$ $\frac{\hat{E}}{f_{A}}\approx 10^{8}$
$\check{\ell}$ $=$ $\frac{\check{E}}{f_{A}}\approx 10^{-8},$ (C1)
where the value of the axion decay constant, $f_{A}\approx 10^{10}~{\rm MeV}$, given in Section 2, ensures the condition $\hat{\ell}\check{\ell}=1$, which has been suggested previously on geometric grounds (R. Penrose, 2016).
With this conformal interpretation, the cyclic cosmology matching condition, given in Eq. (29), simply expresses the conformal rescaling of the Ricci curvature scalar at the cross-over, $R$, to its value at future infinity, $\hat{R}$, and at the Big Bang, $\check{R}$, given by (R. Penrose & W. Rindler, 1986, pg. 123)
$\hat{R}\hat{\ell}^{2}=R=\check{R}\check{\ell}^{2}.$ (C2)
To make the connection, we use the field equations of general relativity to express the value of the Ricci curvature scalar at future infinity, $\hat{R}$, in terms of the cosmological constant, $\Lambda$, and the value at the Big Bang, $\check{R}$, in terms of the trace of the baryon stress-energy tensor at the Big Bang (R. Penrose & W. Rindler, 1986, pg. 20)
$\hat{R}$ $=$ $4\Lambda$
$\check{R}$ $=$ $\frac{8\pi G}{c^{4}}\left(\frac{T_{\pi}}{T_{\gamma}}\right)^{3}\left(\frac{1+\epsilon}{1+\delta}\right)u_{B},$ (C3)
where $u_{B}$ is the energy density of baryons in the present universe, $T_{\pi}$ is the Big Bang temperature, and $T_{\gamma}$ is the temperature of the cosmic microwave background. Combining Eq. (C), (C2), and (C), we arrive at the cyclic cosmology matching condition
$4\Omega_{\Lambda}h^{2}\left(\frac{\hat{E}}{f_{A}}\right)^{2}=\Omega_{B}h^{2}\left(\frac{T_{\pi}}{T_{\gamma}}\right)^{3}\left(\frac{\check{E}}{f_{A}}\right)^{2}\left(\frac{1+\epsilon}{1+\delta}\right),$ (C4)
where $\Omega_{\Lambda}h^{2}=(\Lambda c^{2}/3)(h/H_{0})^{2}$ is the cosmological parameter for $\Lambda$, and $\Omega_{B}h^{2}=(8\pi Gu_{B}/3c^{2})(h/H_{0})^{2}$ is that for $u_{B}$.
We estimate the small dimensionless quantity, $\epsilon=(4m_{\alpha}/m_{p}-1)Y_{P}=1.876~(34)\times 10^{-3}~[18~{\rm per~mille}]$, using the best present value for the ratio, $m_{\alpha}/m_{p}=3.972~599~690~252~(70)~[18~{\rm parts~per~trillion}]$ (P. J. Mohr et al., 2025), of the mass of the helium nucleus to the proton mass, and using the astrophysical determination of the primordial helium abundance within the present physical picture, $Y_{P}=Y_{\rm HB}=0.272~(5)~[18~{\rm per~mille}]$ (See Section 8). The exact expression for the relativistic correction, $1+\delta=m_{p}\eta/(m_{p}e^{-x_{p}}x_{p}J(x_{p})+m_{n}e^{-x_{n}}x_{n}J(x_{n}))$, uses the nucleon mass-to-temperature ratios at the Big Bang, $x_{N}=m_{N}c^{2}/kT_{\pi}$, for $N=p$ and $n$, with $kT_{\pi}$ from Eq. (3). The numerically exact integral, $J(x_{N})$, is defined in Eq. (25), while the integral, $I(x_{N})=(1/2!\zeta(3))\int_{0}^{\infty}dxe^{-x}(x+x_{N})\sqrt{x(x+2x_{N})}$, determines the baryon-to-photon number density ratio, $\eta=e^{-x_{p}}I(x_{p})+e^{-x_{n}}I(x_{n})=1.290~113~(5)\times 10^{-9}~[{\rm 4~ppm}].$
Appendix D 0.5 eV QCD Axion Dark Energy
The derivation of the expression for the cosmological parameter of the cosmological constant, $\Omega_{\Lambda}h^{2}$, in Eq. (27), begins with the physical picture for the nature of dark energy presented in Section 2, but evaluated now at future infinity instead of the Big Bang:
$\Omega_{\Lambda}h^{2}$ $=$ $\frac{8\pi G}{3c^{2}}\left(\frac{h}{H_{0}}\right)^{2}\hat{U}_{\Lambda}\hat{n}_{A}$ (D1)
$=$ $\frac{8\pi}{3}\frac{\hbar c^{3}}{E_{\rm Pl}^{2}}\left(\frac{h_{0}}{H}\right)^{2}\left(N_{U}f_{A}\right)\left(N_{n}\left(\frac{\check{E}}{\hbar c}\right)^{3}\left(\frac{\check{E}}{\hat{E}}\right)^{3}\right)$
$=$ $\frac{8\pi}{3}N_{U}N_{n}\left(\frac{\check{E}^{9}}{f_{A}^{5}E_{\rm Pl}^{2}}\right)\left(\frac{m_{A}c^{2}}{m_{A}c^{2}}\right)^{5}\left(\frac{h}{H_{0}}\frac{1}{\hbar}\right)^{2}$
$=$ $\frac{8\pi}{3}N_{U}N_{n}x_{k}^{9}y_{T}^{10}\frac{\left(m_{A}c^{2}\right)^{5}}{kT_{\pi}E_{\rm Pl}^{2}}\left(\frac{h}{H_{0}}\frac{1}{\hbar}\right)^{2}$
$=$ $\frac{16\zeta(3)x_{k}^{6}y_{T}^{10}}{\pi^{3}}\frac{\left(m_{A}c^{2}\right)^{5}}{kT_{\pi}E_{\rm Pl}^{2}}\left(\frac{h}{H_{0}}\frac{1}{\hbar}\right)^{2},$
where $\hat{U}_{\Lambda}=N_{U}\hat{E}^{3}/E_{\rm Pl}^{2}=N_{U}f_{A}$ is the dark energy inside each axion at future infinity, and $\hat{n}_{A}=(\check{\ell}/\hat{\ell})^{3}(T_{\pi}/T_{\gamma})^{3}n_{A}=N_{n}(\check{E}/\hbar c)^{3}(\check{E}/\hat{E})^{3}$ is the axion number density at future infinity. See Appendix A for the derivation of the dimensionless constants, $N_{U}=1/(2\pi^{2})$ and $N_{n}=12\zeta(3)/(\pi^{2}x_{k}^{3})$, but see Appendix C for the conformal scale factors, $\hat{\ell}=\hat{E}/f_{A}$ and $\check{\ell}=\check{E}/f_{A}$. The see-saw relation, $\check{E}\hat{E}=f_{A}^{2}$ (See Section 2), was used in the third line, while the relations, $\check{E}=x_{k}kT_{\pi}$, and, $m_{A}c^{2}f_{A}=\sqrt{\chi_{\rm QCD}}=(kT_{\pi}/y_{T})^{2}$, were used in the fourth line.
Appendix E 0.5 eV QCD Axion-Photon Coupling Strength
The strength of the axion-photon coupling, $g_{A\gamma\gamma}$, that enters the Lagrangian density, $\mathcal{L}_{A\gamma}=g_{A\gamma\gamma}\phi_{A}{\bf E}\cdot{\bf B}$, for the interaction between the axion field, $\phi_{A}$, and the dot product, ${\bf E}\cdot{\bf B}$, of the electric field, $\bf E$}, and the magnetic field, $\bf B$}, is determined by the decay constant of the axion, $f_{A}$, and by the dimensionless constant, $C_{A\gamma}=\mathcal{O}(1)$, whose value is of order one (F. Wilczek, 1978; S. Weinberg, 1978)
$g_{A\gamma\gamma}$ $=$ $C_{A\gamma}\frac{\alpha}{2\pi}\frac{1}{f_{A}}$ (E1)
$=$ $1.048~854~09~(20)\times 10^{-13}~{\rm MeV}^{-1}~C_{A\gamma}~[0.19~{\rm ppm}],$
where $\alpha=e^{2}/(\hbar c)=1/(137.035~999~18~(2))$ (P. J. Mohr et al., 2025) [1.5 parts per billion] is the best present estimate for the dimensionless electromagnetic “fine-structure” coupling constant, and where we have used, in the second line, the relation for the QCD axion decay constant, $f_{A}=\sqrt{\chi_{\rm QCD}}/(m_{A}c^{2})$ (S. Weinberg, 1978; F. Wilczek, 1978), together with the values for the axion rest-mass energy, $m_{A}c^{2}$, given in Eq. (6), and the QCD topological susceptibility, $\chi_{\rm QCD}$, given by Eq. (3) (See Appendix B for the estimate of the uncertainty in the value of $f_{A}$). To estimate the value of $C_{A\gamma}$, we use its relation with the light-quark mass ratio, $m_{u}/m_{d}=0.485~(19)$ (Z. Fodor et al., 2016), and with the electromagnetic-to-strong axial anomaly ratio, $E/N=8/3$, (R. D. Peccei, 1989; M. Gorghetto & G. Villadoro, 2019),
$C_{A\gamma}$ $=$ $\frac{E}{N}-\frac{2}{3}\frac{4m_{d}+m_{u}}{m_{d}+m_{u}}$ (E2)
$=$ $0.65~(2)~[26~{\rm per~mille}],$
where the value of the anomaly ratio is set by the fact that all twelve flavors of “normal” quarks and leptons within the standard model of particle physics transform with the same axial charge, $Q_{A}=+1$, under the continuous axial $U(1)_{A}$ phase symmetry, $U_{A}(\alpha)=\exp(iQ_{A}\alpha)$, while the opposite axial charge, $Q_{A}=-1$, holds for “mirror” quarks and leptons which have the same correlation between chirality and helicity as “normal” antimatter (Right-handed mirror matter feels the weak nuclear force but left-handed mirror matter does not feel it; See L. M. Lederman & D. N. Schramm (1989), pg. 121–123): The 0.5 eV QCD axion dark matter, $A_{M}$, with a given flavor, $M$, is the unique quantum state formed by superposing the bound state of mirror matter and normal antimatter with the bound state of normal matter and mirror antimatter, such that the particle and the antiparticle with flavor $M$ inside the bound states have the same helicity and chirality, and such that the resulting quantum state for the axion, and hence the axion field, $\phi_{A}$, has the same discrete space-time symmetry properties as the dot product, ${\bf E}\cdot{\bf B}$, ensuring that the Lagrangian density, $\mathcal{L}_{A\gamma}$, is invariant under these transformations (charge-conjugation, space-inversion, and time-reversal; See L. M. Lederman & D. N. Schramm (1989), pg. 121–123). Finally, we combine Eq. (E1) and Eq. (E2) to find the value for the axion-photon coupling strength given in Eq. (7).